OpenStudy (anonymous):

3x+y=24 and -3x+5y=-6 Working out please!

OpenStudy (anonymous):

3x+y=24------(a) -3x+5y=-6----(b) adding both equations u get 6y=18 => y=3 so x= (24-y)/3 ....from eq (a) x=(24-3)/3=21/7=3 therefore x=y=3

OpenStudy (anonymous):

equation (a) + equation (b) gives 3x+y=24------(a) -3x+5y=-6----(b) ___________________ 0 + 6y = 24 - 6 6y = 18 y= 18/6 y=3 put this y in equation (a) to solve for x 3x+3 = 24 3x = 21 x=21/3 = 7 so x=7; y=3